# HiPhO: How Far Are (M)LLMs from Humans in the Latest High School Physics Olympiad Benchmark?

### v3

@@ 1 Introduction
In our large-scale evaluation of 30 state-of-the-art (M)LLMs, we observe a clear performance hierarchy. Closed-source reasoning MLLMs dominate the medal table, winning 6–12 gold medals in 13 Olympiads. However, even the strongest models, such as Gemini-2.5-Pro and GPT-5, still fall short of the very best human contestants, especially in challenging exams like IPhO and EuPhO. In contrast, open-source chat MLLMs failed to secure any gold medals, with most scoring only at or below the bronze threshold. More encouragingly, several open-source reasoning (M)LLMs, including Intern-S1 and DeepSeek-R1, each achieved 4–6 gold medals, particularly in relatively easier exams such as F=MA. Taken together, these results reveal the strong but still non-parity performance of closed-source models, the evident limitations of open-source chat MLLMs, and the promising trajectory of open-source reasoning (M)LLMs in advancing physics problem solving.
@@ 4 Evaluation Method
Key Improvements in Evaluation. Our evaluation introduces three major enhancements. (1) Step-level scoring: Most datasets assess only the final answer (Table 1), awarding zero credit if it is incorrect. We incorporate step-level scoring to recognize partial correctness. (2) Marking-scheme alignment: Unlike prior step-level methods (e.g., OlympicArena generates steps with GPT-4), we are the first to extract steps strictly from the official marking scheme, awarding points based on explicit marking criteria for greater rigor and consistency. (3) Stronger grader: We employ Gemini-2.5-Flash as the grading model. Previous works (Xu et al., 2025; Feng et al., 2025) often rely on GPT-4o, whose competition scores are relatively low, making it less suitable for evaluating Olympiad-level problems. In our experiments, Gemini-2.5-Flash produced grades more closely aligned with human experts (see Appendix B.4).
@@
[…]
@@
DeepSeek-R1 18.5 24.6 25.4 10.8 21.4 26.3 20.5 22.4 20.3 48.5 63.8 18.3 18.5  6 5 2
Qwen3-32B 15.7 19.3 23.9 9.8 21.2 28.9 24.1 26.8 17.9 50.1 58.7 18.9 16.6  6 5 2
Qwen3-235B-A22B 17.8 23.8 26.0 9.2 21.5 28.4 31.1 26.0 18.9 48.9 40.4 18.4 17.4  6 4 3
@@
Qwen3-30B-A3B 13.6 15.4 22.7 9.8 16.5 24.7 21.7 22.8 17.8 40.3 44.6 15.5 15.0  2 9 2
DeepSeek-V3 13.6 16.4 22.1 7.1 17.2 21.1 17.3 23.5 16.9 36.7 44.8 14.1 15.6  1 8 4
Qwen3-8B 10.6 12.7 11.5 7.1 11.9 20.1 17.3 21.2 11.5 15.0 22.5 10.8 10.0  0 2 8
@@ 5.1 Main Results
Can open-source models catch up with closed-source champions? While closed-source reasoning models still dominate, recent progress in the open-source community is noteworthy. Intern-S1 distinguished itself as the only open-source MLLM to win gold, obtaining four medals. Even more impressively, open-source LLMs including DeepSeek-R1, Qwen3-32B, and Qwen3-235B-A22B all achieved six gold medals. These advances reveal both the enduring gap with closed-source leaders and the emerging potential of open-source research to narrow it.
@@ 5.2 Can SOTA (M)LLMs Compete with Human Medalists in Physics Olympiads?
Mechanics-focused Olympiads demonstrate the potential of open-source (M)LLMs to rival human gold medalists. In PanMechanics-2025, which targets younger high-school students and places strong emphasis on classical mechanics, open-source (M)LLMs not only reached the gold threshold but in some cases even surpassed top human scores. For example, Kimi-K2-Instruct achieved 65.9 points compared to the top human score of 62.0. These results suggest that models are already approaching near-expert competence in mechanics, where reasoning is more deterministic and better aligned with training priors, while continuing to struggle in Olympiads that include optics problems, such as EuPhO-2025 and PanPhO-2025.

### v4

@@ 1 Introduction
In our large-scale evaluation of 30 state-of-the-art (M)LLMs, we observe a clear performance hierarchy. Closed-source reasoning MLLMs dominate the medal table, winning 6–12 gold medals in 13 Olympiads. However, even the strongest models, such as Gemini-2.5-Pro and GPT-5, still fall short of the very best human contestants, especially in challenging exams like IPhO and EuPhO. In contrast, open-source chat MLLMs failed to secure any gold medals, with most scoring only at or below the bronze threshold. More encouragingly, several open-source reasoning (M)LLMs, including Intern-S1 and DeepSeek-R1, each achieved 4–8 gold medals, particularly in relatively easier exams such as F=MA. Taken together, these results reveal the strong but still non-parity performance of closed-source models, the evident limitations of open-source chat MLLMs, and the promising trajectory of open-source reasoning (M)LLMs in advancing physics problem solving.
@@ 4 Evaluation Method
Key Improvements in Evaluation. Our evaluation introduces three major enhancements. (1) Step-level scoring: Most datasets assess only the final answer (Table 1), awarding zero credit if it is incorrect. We incorporate step-level scoring to recognize partial correctness. (2) Marking-scheme alignment: Unlike prior step-level methods (e.g., OlympicArena generates steps with GPT-4), we are the first to extract steps strictly from the official marking scheme, awarding points based on explicit marking criteria for greater rigor and consistency. (3) Stronger grader: We employ Gemini-2.5-Flash as the grading model. Previous works (Xu et al., 2025; Feng et al., 2025) often rely on GPT-4o, whose competition scores are relatively low, making it less suitable for evaluating Olympiad-level problems. Compared to GPT-4o, Gemini-2.5-Flash provides step-level scores more closely aligned with human experts, with under 1-point differences in the examples (see Appendix B.4).
@@
Gold Medal 19.7 20.8 23.3 16.5 20.4 28.6 26.5 41.5 52.0 52.0 51.0 15.0 14.0 5
Silver Medal 12.1 11.1 18.7 9.8 14.2 20.1 19.4 28.5 37.5 36.0 26.0 11.0 12.0 5
Bronze Medal 7.2 3.6 13.1 5.8 8.9 15.2 13.5 14.5 16.0 20.0 12.0 9.0 10.0 5
@@
DeepSeek-R1 18.5 24.6 25.4 10.8 21.4 26.3 20.5 42.2 47.4 65.4 72.5 18.3 18.5  8 5 0
Qwen3-235B-A22B 17.8 23.8 26.0 9.2 21.5 28.4 31.1 42.3 49.1 63.1 44.6 18.4 17.4  8 4 1
Qwen3-32B 15.7 19.3 23.9 9.8 21.2 28.9 24.1 36.6 41.8 67.0 59.2 18.9 16.6  7 6 0
@@
Qwen3-30B-A3B 13.6 15.4 22.7 9.8 16.5 24.7 21.7 31.9 39.5 49.9 45.0 15.5 15.0  2 11 0
DeepSeek-V3 13.6 16.4 22.1 7.1 17.2 21.1 17.3 37.2 35.0 48.4 46.5 14.1 15.6  1 9 3
Qwen3-8B 10.6 12.7 11.5 7.1 11.9 20.1 17.3 26.3 22.3 21.8 22.8 10.8 10.0  0 2 10
@@ 5.1 Main Results
Can open-source models catch up with closed-source champions? While closed-source reasoning models still dominate, recent progress in the open-source community is noteworthy. Intern-S1 distinguished itself as the only open-source MLLM to win gold, obtaining four medals. Even more impressively, open-source LLMs including DeepSeek-R1 and Qwen3-235B-A22B both achieved eight gold medals. These advances reveal both the enduring gap with closed-source leaders and the emerging potential of open-source research to narrow it.
@@ 5.2 Can SOTA (M)LLMs Compete with Human Medalists in Physics Olympiads?
Mechanics-focused Olympiads demonstrate the potential of open-source (M)LLMs to rival human gold medalists. In PanMechanics-2025, which targets younger high-school students and places strong emphasis on classical mechanics, open-source (M)LLMs not only reached the gold threshold but in some cases even surpassed top human scores. For example, Kimi-K2-Instruct achieved 65.9 points compared to the top human score of 62.0. These results suggest that models are already approaching near-expert competence in mechanics, where reasoning is more deterministic and better aligned with training priors, while continuing to struggle in Olympiads that include optics problems, such as EuPhO-2025 and PanPhO-2024.